OpenStudy (anonymous):

An electron travels 1.92 m in 4.47 × 10−8 s. How fast is it traveling? Answer in units of m/s

OpenStudy (anonymous):

1.92/(4.47*E-8) = ... [m/s]

OpenStudy (anonymous):

You're seeking units of m/s and have been supplied meters and seconds so: \[\frac{1.92m}{4.47*10^{-8}s}\]

OpenStudy (anonymous):

Plugging in a calculator you get \[42953020.13 \frac{m}{s}\] Odds are you need to supply this in scientific notation: \[4.30*10^7\frac{m}{s}\]

OpenStudy (anonymous):

Turner’s treadmill starts with a velocity of −2.7 m/s and speeds up at regular intervals during a half-hour workout.After 33 min, the treadmill has a velocity of −7.1 m/s. What is the average acceleration of the treadmill during this period? Answer in units of m/s2